# Low-Latency ML Inference by Grouping Correlated Data Objects and Computation

### v1

@@ 4.6. Results
We now report results for experiments with just one client. We varied the number of shards for MOT (1, 3), PRED (1, 3, 5) and CD (1, 3, 5). The number of nodes in each shard was 1. Figure 4 shows a box plot of the E2E latency (in milliseconds) for different layouts and placement strategies. We show only results for client gates3 for the sake of saving space. Results for other clients followed the same pattern. It is possible to see that correlation grouping significantly decreased both median and 75th percentile latencies, except for layout 1/1/1, in which the grouping mechanism had no effect since there was only 1 shard per step. Furthermore, increasing the number of shards did not improve latency with the random placement strategy. In fact, both the median and percentile latencies slightly increased as more shards were added: fetching overhead increased, due to more cache misses.
@@
We now increase the workload to three simultaneous clients. Figure 5 shows the E2E latency (in milliseconds) for three simultaneous clients (little3 + hyang5 + gates3), employing different layouts and placement strategies. Latency was significantly lower and more consistent as the deployment scaled out when employing correlation grouping.
@@
Figure 5. E2E latency for three clients on Cascade.
In order to further highlight the benefit of collocation, we disabled the object caching performed by each step of our application. Without caching, steps always have to fetch objects from the K/V store. In this context, the main difference between the random placement and correlation grouping strategies is that correlation grouping guarantees that objects being fetched are always stored locally, i.e. in the same Cascade node (and shard) where the computation is running. In the case of random placement, there is no guarantee: objects are placed randomly across shards.
Figure 6 shows the results comparing the two placement strategies with and without the application-level caching, for a workload of three simultaneous clients, and 3/5/5 shards (MOT/PRED/CD steps, respectively), one node per shard. Due to the zero-copy design of Cascade, E2E latency was the same with or without caching when employing correlation grouping, since memory copies are minimized and there is no serialization overhead when an UDL makes a get request to the same node where it is running: keeping objects in the application memory or fetching them locally from the K/V store incurs virtually the same cost. However, in the case of random placement, disabling caching significantly increased latency and reduced throughput. The median E2E latency was off-scale at more than 58 seconds, hence the bar for this case is replaced with an arrow. Throughput was on average 6.7 FPS, thus the pipeline was not able to handle the 7.5 FPS sent by clients collectively.
@@
We also evaluated the impact of replication on the E2E latency of the application. Different layouts with more than one node per shard were employed for this evaluation. On Cascade, when a shard has more than one node, any object stored in that shard is replicated to all nodes in the shard, before any computation is triggered. As a result, any computation triggered in such a shard will have local access to the objects in that shard, even if the previous computation (e.g. state of the previous frame) was performed in another node. This behavior can be seen as a form of prefetching, since nodes store objects that will be used in the future before the corresponding request arrives. However it incurs extra latency since the corresponding computation is only triggered after the replication is completed, while a more sophisticated prefetching feature would not prevent the computation to start while prefetching is done in other nodes.
Figure 7 shows results for different layouts. The first group of bars is presented as reference, and show the E2E latency for 3/5/5 shards, with one node per shard (no replication). The next two bars present results for 1/1/1 shards, with different number of nodes per shard (3 and 3/5/5), so there is no difference between random placement and correlation grouping. The last layout has 1/3/3 shards, with 2 nodes per shard, configuring a compromise between having many shards with just one node each and a single shard with many nodes. Results show that replication reduced latency compared to the baseline (first bar). However, employing correlation grouping and multiple shards still results in better E2E latency. We argue that these results show the potential of further investigating the integration of correlation grouping with prefetching and scheduling.
@@ 4.7. Insights
Correlation grouping significantly reduces the latency of the RCP application on Cascade, compared to other placement options Cascade is able to support. Additional benefits include better scalability than with random placement, where adding nodes sometimes degraded performance. Performance was steady across varied workloads, suggesting that correlation grouping improves hardware utilization. Changes to the application were primarily to register regular expressions (Table 1 during initialization, as illustrated in Figure 3). Support for correlation grouping required minor changes to Cascade itself, yet enabled the system to use correlation keys to place objects and route requests. Our results on replication indicate that there is potential to improve latency even further by integrating correlation grouping with prefetching.
5. RCP Deployment on a Public Cloud
A natural question to ask is whether our claim that existing SP and ML serving platforms are inadequate is correct. To address this, we now describe a second deployment of RCP on Microsoft Azure Cloud. Granted, Azure is a public cloud whereas Cascade primarily targets a private cluster, but we believe this is a fair comparison because it genuinely represents a state-of-the-art alternative.
The goal of the section is to show evidence that: (i) data access patterns of applications such as RCP indeed result in extra fetching overheads; and (ii) although it is possible to overcome these overheads, it comes at the cost of a high coupling between application and deployment, since there is no unified mechanism available in modern platforms such as the proposed correlation grouping.
@@ 5.1. Deployment on Microsoft Azure Cloud
The RCP pipeline was implemented on Stream Analytics (SA) – Azure’s SP platform. We reused the source-code of each of the 3 steps (MOT, PRED, and CD), with few modifications. Each of the ML models was first configured on Azure Machine Learning (AML) as real-time endpoints, using a web-services interface. AML is elastic: depending on load, endpoints can be backed by worker instance pools of varying size (Microsoft, 2022a). Pipeline stages were connected using Azure Event Hubs (EHs) (Kettner and Geisler, 2022). Clients were deployed in a virtual machine. Frames from all three videos were stored in Azure Blob Storage (Calder et al., 2011), which was also used to store MOT state data. Positions of agents and trajectory predictions are stored using Cosmos DB (Reagan, 2018). The resulting architecture is shown in Figure 8, and is compliant with recommendations from the cloud vendor. The individual ML tasks were properly configured with ample resources (similar to the resources employed previously in the Cascade experiments).
@@
Clients stream video frames to the first SA job in the pipeline at a rate of 2.5 FPS, through an EH. The SA job invokes an MOT endpoint: one of the instances in the endpoint is selected by a load balancer to serve the request. The instance fetches the video frame from Blob storage and performs the inference (we assume the client has already uploaded the frame). The state from the previous frame is also fetched from Blob storage if it is not already cached in the instance’s memory. Results are sent through another EH to the next SA job responsible for invoking PRED endpoints for agent positions. For each invocation, the selected instance of the invoked PRED endpoint stores the new position on Cosmos DB, and also fetches the past positions of the corresponding agent that are not already cached in memory. The CD step is performed in a similar fashion. The final output of the pipeline is sent back to the client running on the virtual machine. End-to-end latency of a frame is the time elapsed from when the frame request was first sent by the client, until the last output from the CD step for that frame was received by the client.
@@
We employed as many partitions in EHs and streaming units in SA as supported by the endpoints (2*number of instances), maximizing parallelism (Microsoft, 2022b). The partition key for MOT requests was the video name, ensuring that frames originating in any single client were processed sequentially. Similarly, the partition key for PRED was the agent identifier, and for CD the frame number. We set the batch size when invoking AML endpoints to 1, a configuration intended to minimize per-frame latency.
@@
5.2. Results: From One to Two Simultaneous Clients
@@
Figure 9 shows a box plot of the E2E latencies (in seconds) for each video and number of PRED/CD instances. The number of instances for PRED and CD endpoints is indicated on the horizontal axis. For gates3, the latency for 1 instance was significantly high and thus the bar was replaced by an arrow. The plot shows that increasing the number of instances from 1 to 3 reduced the median and 75th percentiles. However, no significant benefit was observed when increasing the number of instances to 5 for little3 and hyang5. The reason for this is that 3 instances were enough to significantly parallelize invocations to PRED. Increasing the number of instances even further resulted in more network overhead of fetching agent positions that offset the gain of extra parallelism, since the rate of cache misses increased. With 3 instances, 64ms was spent per frame fetching data in the PRED step for hyang5, while 74ms was spent on average with 5 instances.
Figure 9. End-to-end latency for individual clients.
@@
Figure 10. Latency breakdown of the MOT step.
5.3. Results: Slashing Overheads From MOT
@@
Figure 11. Grouped MOT with 2 and 3 simultaneous clients.
However, we observed that there is still potential to further cut network overheads. As we increased the number of PRED and CD instances, parallelism is increased, but so is the time spent fetching agent positions and trajectory predictions from Cosmos DB. Figure 12 shows the latency breakdown for PRED and CD steps, with three simultaneous clients, three grouped MOT endpoints, and a varying number of PRED and CD instances (3, 5, and 7). The plot shows the average time spent per frame (in ms) fetching the necessary input data and running the corresponding inference. The values shown are a sum of the measured times for all invocations. With 7 PRED/CD instances, for example, the average time spent fetching data per frame in PRED and CD steps was almost 200ms (roughly 100ms for each step).
Figure 12. Latency breakdown of PRED and CD steps.
5.4. Results: Slashing Overheads From PRED and CD
Here we employ the same approach as described above for MOT, leveraging application-specific knowledge to group PRED and CD. We deploy multiple endpoints with one instance each, instead of one endpoint with multiple instances. However, it is not possible to rely any longer on the endpoint load balancer to select which instance will receive each request. Thus, we manually select in the SA jobs code which endpoint to forward each request, by writing the output of the previous step to the corresponding EH. For PRED, this selection was based on the agent identifier, modulo the number of PRED endpoints. For CD, the selection was based on the frame number, modulo the number of CD endpoints.
Figure 13 shows a code excerpt (simplified for the sake of presentation) from the SA job responsible for the MOT endpoint associated with the video little3. In the code, from-little3-client is an EH where the client responsible for the little3 video sends requests. Function mot-little3-endpoint invokes the AML endpoint containing a single instance that performs the MOT inference. There are two queries (lines 9 and 12) that forwards the position of each agent to a specific EH (to-pred-0 or to-pred-1) depending on the agent identifier and the number of PRED endpoints. The EHs to-pred-0 and to-pred-1 will deliver the agents positions to the corresponding SA jobs that will invoke their associated PRED endpoints.
@@ 6  CROSS APPLY GetArrayElements(res) AS r
10WHERE agent_id % num_pred_endpoints = 0
@@
13WHERE agent_id % num_pred_endpoints = 1
Figure 13. Code excerpt from the MOT SA job for little3.
@@
Figure 14. Grouping only MOT and grouping all steps.
@@ 5.5. Insights, Trade-offs, and Limitations
Without our manual grouping, scaling in Azure is straightforward: it is only necessary to increase the number of instances in AML, streaming units in SA, and/or partitions in EHs. AML endpoints will distribute requests across all the available instances automatically according to their load. Furthermore, it is possible to use auto-scaling features, which will scale the deployment in/out according to predefined metrics.
However, when employing the manual grouping strategy, scaling is done by adding or removing endpoints, which requires the application to be reconfigured. In the case of our deployment in Azure, changing endpoints affects client and SA jobs code, and it is also necessary to create new EHs. Load balancing and auto-scaling would need to be implemented by the application itself.
Despite all trade-offs, the benefit of grouping is significant, as also observed in Section 4. For MOT in particular, grouping was essential to support more than one client. Without grouping, the overhead of fetching states caused each MOT request to take more than 400ms, which is the time it takes for each client to send the next frame request. This caused requests to pile up in queues at the first step of the pipeline. In a time-pressured scenario such as the RCP application, the 400ms threshold is of high importance: the ML model employed in the PRED step predicts agents trajectories in the next 4.8 seconds. If it takes too long to process each frame, due to queuing, then the prediction may not be useful any longer. Furthermore, since the application handles a real-time video stream, the MOT state data for each frame will only be ever used once (for the following frame). However, without grouping it is moved twice (upload and download), which can be considered wasteful.
Another clear benefit of grouping is better instance utilization. One could argue that each video/client configures a separate application which could be deployed separately, in a dedicated infrastructure. However, we showed that grouping allows multiple clients to share the same infrastructure while maintaining the same level of performance, which requires less instances to be deployed since they are better utilized. Furthermore, such shared deployment enables new clients to be dynamically added in runtime. However, this would require an extra front-end step in the pipeline that decides how the grouping of a new client should be performed (we’ve done it statically for our experiments).
@@ 6. Related Work
Prior efforts (Lee et al., 2018; Kraft et al., 2020; Park et al., 2022; Ahmed et al., 2019; Paganelli et al., 2023) argue for a white box approach, in which optimizations are employed in compilation or training time, based on specific features such as topology of neural networks. While such approaches can reduce the computational cost of model serving, our work employs a black box approach in which the application only needs to identify data / compute correlations. Although the authors in (Lee et al., 2018) criticize black box approaches, they focus only on caching, buffering and batching in non-pipelined applications. The correlation grouping mechanism can support all of these capabilities but also enables additional optimizations, such as proactive data/computation collocation and prefetching.
@@
Existing systems offer a number of mechanisms to express application-specific knowledge and/or request grouping; these features can be found in Redis hash tag, Azure Cosmos DB, EHs and SA partition key, RDD location preferences, Apache Spark Streaming partitions, and Apache Storm stream grouping. However, as noted in Section 3, such mechanisms: (i) are highly coupled with deployment; (ii) do not collocate data and computation in a unified fashion; (iii) are not able to express all possible correlations between data and computation; and/or (iv) may have a non-negligible computational cost as the deployment scales out.
@@
The Pheromone (Yu et al., 2023) system offers a novel approach that the authors refer to as a data bucket abstraction, in which the outputs of functions are automatically grouped into buckets. Developers can then arrange for downstream functions to use data buckets as inputs. This enables Pheromone to place functions close to where the data is stored. The data bucket abstraction is similar to correlation grouping, however outputs in data buckets are always volatile (once consumed, they are garbage collected). The correlation grouping approach is compatible with data persistence, and we would argue that this gives greater flexibility. For example, some outputs may be needed between separate invocations of a pipeline (e.g. past positions of agents in the PRED step of the RCP application); our approach covers such cases but the data bucket model does not. Furthermore, the authors do not consider more complex workflows in which multiple functions may require access to the same output in different stages, or in which a function requires data located in multiple buckets.
@@
Several works improve data locality by scheduling tasks where input data (or most of it) is located (Nishihara et al., 2017; Moritz et al., 2018; Sreekanti et al., 2020; Yu et al., 2021; Boutin et al., 2014; Crankshaw et al., 2020; Zhang et al., 2017), or by scheduling tasks that interact with each other in the same location (node, VM, or container) (Mahgoub et al., 2022; Kotni et al., 2021). These solutions follow a reactive approach: data is first placed without taking into account their correlations, and then tasks are placed where “most of input data” is located. Due to the data access patterns of latency-sensitive ML pipelines, a proactive approach, such as enabled by correlation grouping, may be necessary. We claim that our proposed mechanism is complementary to scheduler-based solutions.

### v2

@@ 4.6. Results
We now report results for just one client, varying the number of shards for MOT (1, 3), PRED (1, 3, 5) and CD (1, 3, 5). Figure 3 shows a box plot of the E2E latency (in milliseconds) for different layouts and placement strategies. We show only client gates3, results for other clients followed the same pattern. Both median and 75th percentile latencies were significantly reduced by affinity grouping, except for layout 1/1/1: grouping had no effect since there was only 1 shard per step. Adding shards does not help with the random placement strategy. Although increasing from 1/1/1 to 1/3/3 did help, increasing the number of shards even further results in higher median and percentile latencies, as fetching overheads increase due to more cache misses. We now increase the workload. Figure 4 shows the E2E latency for three simultaneous clients (little3 + hyang5 + gates3), employing different layouts and placement strategies. Latency was significantly lower and more consistent as the deployment scaled out when employing affinity grouping.
@@
To further highlight the benefit of collocation, we disabled object caching in our application. Each step then always has to fetch objects from the K/V store. Here, the main difference between random placement and affinity grouping is that the latter guarantees that objects being fetched are always stored locally, in the same Cascade node (and shard) where the task is running. For random placement, there is no such guarantee. Figure 5 shows the results for three clients and 3/5/5 shards. Due to the zero-copy design of Cascade, E2E latency was the same with or without caching for affinity grouping, since memory copies are minimized and there is no serialization overhead when an UDL makes a get request to the same node where it is running: keeping objects in the application memory or fetching them locally from the K/V store incurs virtually the same cost. However, for random placement, disabling caching significantly increased latency and reduced throughput. The median latency was off-scale at more than 58 seconds, hence the bar is replaced with an arrow. Throughput was on average 6.7 FPS, thus the pipeline was not able to handle the 7.5 FPS sent by clients collectively.
@@
We also evaluated the impact of replication on E2E latency. Different layouts with more than one node per shard were employed. In Cascade, when a shard has more than one node, any object stored in that shard is replicated to all nodes in the shard, before any task is triggered. As a result, tasks in such a shard will have local access to the objects in that shard, even if the previous computation (e.g. state of the previous frame) was performed in another node. This behavior can be seen as a form of prefetching, since nodes store objects that will be used in the future before the corresponding request arrives. However it incurs extra latency since the corresponding computation is only triggered after the replication is completed, while a more sophisticated prefetching feature would not prevent the computation to start while prefetching is done in other nodes. Figure 6 shows results for different layouts. As a reference, the first group of bars show the latency for 3/5/5 shards, one node per shard (no replication). The next two bars show results for 1/1/1 shards, with different number of nodes per shard (3 and 3/5/5), so there is no difference between random placement and affinity grouping. The last layout has 1/3/3 shards, two nodes per shard, configuring a compromise between many shards with just one node each and a single shard with many nodes. Results show that replication reduced latency compared to the baseline (first bar). However, employing affinity grouping and multiple shards still results in better latency. We argue that these results show the potential of further investigating the integration of affinity grouping with prefetching and scheduling.
@@ 4.7. Insights
Affinity grouping significantly reduces the latency of the RCP application on Cascade, compared to other placement options. Although the sharding policy in Cascade makes it scalable by design, results show that affinity grouping improved Cascade scalability even further compared to the standard random placement, where adding nodes sometimes degraded performance due to increased cache misses. It is important to note that the scalability of affinity grouping is orthogonal to the scalability of the system as a whole. The scalability of the affinity grouping mechanism itself is discussed in Section 3.
Changes to the application were primarily to register regular expressions (Table 1 during initialization, as illustrated in Listing 1). Support for affinity grouping required minor changes to Cascade itself, yet enabled the system to use affinity keys to place objects and route requests. Our results on replication indicate that there is potential to improve latency even further by integrating affinity grouping with prefetching.
5. Evaluation on Azure Cloud
A natural question to ask is whether our claim that existing SP and AI serving platforms are inadequate is correct. To address this, we now describe a second deployment of RCP on Microsoft Azure Cloud. Granted, Azure is a public cloud whereas Cascade primarily targets a private cluster, but we believe this is a fair comparison because it genuinely represents a state-of-the-art alternative.
The goal of the section is to show evidence that: (i) data access patterns of applications such as RCP indeed result in extra fetching overheads; and (ii) although it is possible to overcome these overheads, it comes at the cost of a high coupling between application and deployment, since there is no unified mechanism available in modern platforms such as the proposed affinity grouping.
@@ 5.1. Deployment on Microsoft Azure Cloud
The RCP pipeline was implemented on Stream Analytics (SA) – Azure’s SP platform. We reused the source-code of each of the 3 steps (MOT, PRED, and CD), with few modifications. Each of the AI models was first configured on Azure Machine Learning (AML) as real-time endpoints, using a web-services interface. AML is elastic: depending on load, endpoints can be backed by worker instance pools of varying size (Microsoft, 2022a). Pipeline stages were connected using Azure Event Hubs (EHs) (Kettner and Geisler, 2022). Clients were deployed in a virtual machine. Frames from all three videos were stored in Azure Blob Storage (Calder et al., 2011), which was also used to store MOT state data. Positions of actors and trajectory predictions are stored using Cosmos DB (Reagan, 2018). The resulting architecture is shown in Figure 7, and is compliant with recommendations from the cloud vendor. The individual AI tasks were properly configured with ample resources (similar to the resources employed previously in the Cascade experiments).
@@
Clients stream frames to the first SA job in the pipeline at 2.5 FPS, through an EH. The SA job invokes an MOT endpoint: one instance in the endpoint is selected by a load balancer to serve the request. The instance fetches the video frame from Blob storage and performs the inference (we assume the client has already uploaded the frame). The state from the previous frame is also fetched from Blob storage if it is not already cached in the instance’s memory. Results are sent through another EH to the next SA job responsible for invoking PRED endpoints for actor positions. For each invocation, the selected instance of the invoked PRED endpoint stores the new position on Cosmos DB, and also fetches the past positions of the corresponding actor that are not already cached in memory. The CD step is performed in a similar fashion. The final output of the pipeline is sent back to the client running on the virtual machine. End-to-end latency of a frame is the time elapsed from when the frame request was first sent by the client, until the last output from the CD step for that frame was received by the client.
@@
We employed as many partitions in EHs and streaming units in SA as supported by the endpoints (2*number of instances), maximizing parallelism (Microsoft, 2022b). The partition key for MOT requests was the video name, ensuring that frames originating in any single client were processed sequentially. Similarly, the partition key for PRED was the actor identifier, and for CD the frame number. We set the batch size when invoking AML endpoints to 1, a configuration intended to minimize per-frame latency.
@@
Figure 8. E2E latency for individual clients.
Figure 9. MOT latency breakdown.
5.2. Increasing Workload
@@
Figure 8 shows a box plot of the E2E latencies (in seconds) for each video and number of PRED/CD instances. The number of instances for PRED and CD endpoints is indicated on the horizontal axis. For gates3, the latency for 1 instance was significantly high and thus the bar was replaced by an arrow. The plot shows that increasing the number of instances from 1 to 3 reduced the median and 75th percentiles. However, no significant benefit was observed when increasing the number of instances to 5 for little3 and hyang5. The reason for this is that 3 instances were enough to significantly parallelize invocations to PRED. Increasing the number of instances even further resulted in more network overhead of fetching actor positions that offset the gain of extra parallelism, since the rate of cache misses increased. With 3 instances, 64ms was spent per frame fetching data in the PRED step for hyang5, while 74ms was spent on average with 5 instances.
@@
Figure 10. Grouped MOT with 2 and 3 clients.
5.3. Slashing Overheads From MOT
@@
However, we observed that there is still potential to further cut network overheads. As we increased the number of PRED and CD instances, parallelism is increased, but so is the time spent fetching actor positions and trajectory predictions from Cosmos DB. Figure 11 shows the latency breakdown for PRED and CD steps, with three simultaneous clients, three grouped MOT endpoints, and a varying number of PRED and CD instances (3, 5, and 7). The plot shows the average time spent per frame (in ms) fetching the necessary input data and running the corresponding inference. The values shown are a sum of the measured times for all invocations. With 7 PRED/CD instances, for example, the average time spent fetching data per frame in PRED and CD steps was almost 200ms (roughly 100ms for each step).
Figure 11. PRED and CD latency breakdown.
5.4. Slashing Overheads From PRED and CD
Here we employ the same approach as described above for MOT, leveraging application-specific knowledge to group PRED and CD. We deploy multiple endpoints with one instance each, instead of one endpoint with multiple instances. However, it is not possible to rely any longer on the endpoint load balancer to select which instance will receive each request. Thus, we manually select in the SA jobs code which endpoint to forward each request, by writing the output of the previous step to the corresponding EH. For PRED, this selection was based on the actor identifier, modulo the number of PRED endpoints. For CD, the selection was based on the frame number, modulo the number of CD endpoints.
Listing 2 shows a code excerpt (simplified for the sake of presentation) from the SA job responsible for the MOT endpoint associated with the video little3. In the code, from-little3-client is an EH where the client responsible for the little3 video sends requests. Function mot-little3-endpoint invokes the AML endpoint containing a single instance that performs the MOT inference. There are two queries (lines 9 and 12) that forwards the position of each actor to a specific EH (to-pred-0 or to-pred-1) depending on the actor identifier and the number of PRED endpoints. The EHs to-pred-0 and to-pred-1 will deliver the actors positions to the corresponding SA jobs that will invoke their associated PRED endpoints.
Listing 2: Code excerpt for the MOT SA job.
@@ 6 CROSS APPLY GetArrayElements(res) AS r
10WHERE actor_id % num_pred_endpoints = 0
@@
13WHERE actor_id % num_pred_endpoints = 1
@@
Figure 12. Grouping only MOT vs grouping all steps.
@@ 5.5. Insights, Trade-offs, and Limitations
Without our manual grouping, scaling in Azure is straightforward: it suffices to increase the number of AML instances, SA streaming units and/or EH partitions. AML endpoints will distribute requests across all the available instances automatically according to their load. Furthermore, the pattern lends itself to autoscaling. In contrast, with manual grouping scaling entails adding or removing endpoints, which requires that the application be reconfigured. In the case of Azure, changing endpoints requires code changes to client and SA jobs and the creation of new EHs, and the application itself would have to do load-balancing/auto-scaling.
The benefit of grouping is significant, echoing our observation in Section 4. For MOT, grouping was essential to support more than one client. Without grouping, the overhead of fetching states inflates runtimes to more than 400ms, meaning that the next frame arrives before the current frame is processed (at 2.5 FPS). Requests will pile up at the first pipeline stage. For a time-pressured scenario such as the RCP application, the 400ms threshold is of high importance: the AI model employed in the PRED step predicts actors trajectories in the next 4.8 seconds. If it takes too long to process each frame, due to queuing, it will be too late to act. Furthermore, since the application handles a real-time video stream, the MOT state data for each frame will only be used once (by the subsequent frame), yet without grouping is moved twice (upload and download).
@@ 6. Related Work
Prior efforts (Lee et al., 2018; Kraft et al., 2020; Park et al., 2022; Ahmed et al., 2019; Paganelli et al., 2023) argue for a white box approach, in which optimizations are employed in compilation or training time, based on specific features such as topology of neural networks. While such approaches can reduce the computational cost of model serving, our work employs a black box approach in which the application only needs to identify data/compute correlations. Although the authors in (Lee et al., 2018) criticize black box approaches, they focus only on caching, buffering and batching in non-pipelined applications. The affinity grouping mechanism can support all of these capabilities but also enables additional optimizations, such as proactive data/computation collocation and prefetching.
@@
Existing systems offer a number of mechanisms to express application-specific knowledge and/or request grouping; these features can be found in Redis hash tag, Azure Cosmos DB, EHs and SA partition key, RDD location preferences, Apache Spark Streaming partitions, and Apache Storm stream grouping. However, as noted in Section 3, such mechanisms: (i) are highly coupled with deployment; (ii) do not collocate data and computation in a unified fashion; (iii) are not able to express all possible correlations between data and computation; and/or (iv) may have a non-negligible computational cost as the deployment scales out. We compare affinity grouping with these other approaches in more detail in Section 3.2.
@@
The Pheromone (Yu et al., 2023) system offers a data bucket abstraction, in which the outputs of functions are automatically grouped into buckets. Developers can then arrange for downstream functions to use data buckets as inputs. This enables Pheromone to place functions close to where the data is stored. The data bucket abstraction is similar to affinity grouping, however outputs in data buckets are always volatile (once consumed, they are garbage collected). The affinity grouping approach is compatible with data persistence, and we would argue that this gives greater flexibility. Furthermore, the authors do not consider more complex workflows in which multiple functions may require access to the same output in different stages, or in which a function requires data from multiple buckets.
@@
Several works improve data locality by scheduling tasks where input data (or most of it) is located (Nishihara et al., 2017; Moritz et al., 2018; Sreekanti et al., 2020; Yu et al., 2021; Boutin et al., 2014; Crankshaw et al., 2020; Zhang et al., 2017), or by scheduling tasks that interact with each other in the same location (node, VM, or container) (Mahgoub et al., 2022; Kotni et al., 2021). These solutions follow a reactive approach: data is first placed without taking into account their correlations, and then tasks are placed where “most of input data” is located. Due to the data access patterns of latency-sensitive AI pipelines, a proactive approach, such as enabled by affinity grouping, may be necessary. We claim that our proposed mechanism is complementary to scheduler-based solutions.
Finally, data collocation is widely explored in literature and thus is not a novel concept per se (Mazumdar et al., 2019). However, we argue that our novelty lies in providing developers with an easy-to-use mechanism that requires no knowledge of deployment/environment details, while still achieving effective collocation. The experiments in Appendix 5 are evidence of the limitation of modern platforms and the potential to improve latency by applying collocation principles. Nevertheless, achieving such collocation is challenging for developers and pose trade-offs, hence the need for a better mechanism.
7. Beyond RCP and Cascade
In this work we employed the RCP application as a running example, and implemented affinity grouping only on the Cascade platform. A natural question that arises is whether affinity grouping can be generalized to other applications and/or AI hosting platforms. To address such question, we discuss next how the affinity grouping mechanism could be implemented in Azure, as well as other emerging applications where affinity grouping can have a great impact.
7.1. Affinity grouping in Azure
In Section 4 we described an implementation on a platform based on a K/V store. Few modifications would be required to implement our proposals in Azure. For example, if Azure’s storage solutions and AML operated in a consistent fashion, data stored in Cosmos DB or Blob storage with a certain affinity key could be prefetched by physical servers running the AML instance handling requests with that affinity key. AML endpoints with multiple instances should ideally route requests with the same affinity key to the same instance. If the load balancer will add instances to reduce load on a hot-spot, it could prefetch correlated data to pre-warm those instances. Azure has many caching components; these could be extended to load or evict object sets, applying the identical policy objects sharing a single affinity key.
7.2. Emerging applications
In reinforcement learning (RL) (Sheng et al., 2025), an ML system might run for hours or days repeatedly accessing the same data objects, but not always using the same nodes for running tasks. A developer of such system knows which data objects are accessed by each task, and therefore can map affinity keys to tasks and data objects. Such mapping allows a node to fetch all needed objects for a task invocation (they all share the same affinity key) at once and in parallel, and cache them as a set (or evict them as a set). Such objects may be a set of cached K/V pairs, corresponding to the activation state of the training NN, needed by a Low Rank Adapter (LoRA). Fetching one by one incurs significantly higher delays, because it doesn’t leverage the parallelism of the network, and caching systems lack semantic knowledge of the application, which is clear for the developer and can be easily encoded through affinity grouping.
The same rationale above also applies in the training of modern transformer-based models (e.g. LLMs). In such context, it is common practice to cache collections of Query ($Q$), Key ($K$), Value ($V$) tuples, which are a snapshot of the DNN state of some layer $L$ at step $T$, triggered by input query Q. Any given ($Q,K,V$) tuple is updated by only one neuron in layer $L$, but the iterated (“auto-regressive”) training algorithm requires that at each step every neuron in layer $L$ read the full set of layer $L$ tuples from the prior step, adapt its own weights (compute a gradient), apply the gradient to the layers around it (forward and back propagation), and then update its (Q, K, V) tuple. By hashing $Q$, $L$ and $T$ we can identify an affinity group: the entire group of tuples for this query at this time step for layer $L$ can be cached (or evicted) jointly. Updates will stream from individual neurons to the tuples they own, and as each step finishes, the neurons active in the next step will simply read the cached prior values.
In contrast, suppose we viewed this affinity grouping as a key in a standard K/V store like Reddis. Clearly, we could form a single object collecting the set of ($Q,K,V$) tuples for layer $L$ at step $T$. Once serialized to a byte vector, we would have a K/V object that can be stored in a sharded key-value storage service. But because this one big object has contributions from all neurons in $L$, we would either need to run AllReduce simply to compute it, or pick a K/V store that supports coherent in-place updates and have each neuron read the working value, update it to add its own ($Q,K,V$) tuple, and write the result back. After one update per neuron, the version of the object would be complete for step $T$, and step $T+1$ could run. The issue here is that all the neurons in layer $L$ contribute to this object, and because they run concurrently, all contend to read, modify, and rewrite the object when creating it. That will take time: one layer can have many neurons, so these two alternative approaches (AllReduce or one big object holding the collection) would be very slow in comparison to an affinity-based solution, where each node only reads tuples from its own cache.